OpenStudy (anonymous):

find f(3)if f(x)=(x-1)/(x+2)

OpenStudy (anonymous):

i believe its 2/5 but im not sure ..

OpenStudy (mertsj):

f(3) = (3-1)(3+2)= 2/5